### v1

# Proactive and Reactive Constraint Programming for Stochastic Project Scheduling with Maximal Time-Lags

Abstract
This study investigates scheduling strategies for the stochastic resource-constrained project scheduling problem with maximal time lags (SRCPSP/max)). Recent advances in Constraint Programming (CP) and Temporal Networks have re-invoked interest in evaluating the advantages and drawbacks of various proactive and reactive scheduling methods. First, we present a new, CP-based fully proactive method. Second, we show how a reactive approach can be constructed using an online rescheduling procedure. A third contribution is based on partial order schedules and uses Simple Temporal Networks with Uncertainty (STNUs). Our statistical analysis shows that the STNU-based algorithm performs best in terms of solution quality, while also showing good relative offline and online computation time.
1 Introduction
In real-world scheduling applications, durations of activities are often stochastic, for example, due to the inherent stochastic nature of processes in biomanufacturing. At the same time, hard constraints must be satisfied: e.g. once fermentation starts, a cooling procedure must start at least (minimal time lag) 10 and at most (maximal time lag) 30 minutes later. The combination of maximal time lags and stochastic durations is especially tricky: a delay in duration can cause a violation of a maximal time lag when a resource becomes available later than expected. Such constraints are reflected in the Stochastic Resource-Constrained Project Scheduling Problem with Time Lags (SRCPSP/max). This problem has been an important focus of research due to its practical relevance and the computational challenge it presents, as finding a feasible solution is NP-hard [BMR88].
Broadly speaking, there are two main schools of thought regarding solution approaches for stochastic scheduling in the literature: 1) proactive scheduling and 2) reactive scheduling. The main goal of proactive scheduling is to find a robust schedule offline, whereas reactive approaches adapt to uncertainties online. Proactive and reactive approaches can be considered as opposite ends of a spectrum. Both in practice and literature, it is often observed that methods are hybrid, such as earlier work on the SRCPSP/max.
Hybrid approaches appear for example in the form of a partial order schedule (POS), which is a temporally flexible schedule in which resource feasibility is guaranteed [PSCO04]. The state-of-the-art POS approach for SRCPSP/max is the algorithm $\mathsf{BACCHUS}$ [FVL16], although the comparison provided by the authors themselves shows that their earlier proactive method $\mathsf{SORU}$-$\mathsf{H}$ [VFL16] performs better. Partial order schedules are often complemented with a temporal network [LM09], in which time points (nodes) are modeled together with temporal constraints (edges). Recent advances in temporal networks with uncertainties [HP24] pave the way for improvements in POS approaches for SRCPSP/max.
State-of-the-art methods like $\mathsf{BACCHUS}$ and $\mathsf{SORU}$-$\mathsf{H}$ use Mixed Integer Programming (MIP), whereas Constraint Programming (CP), especially with interval variables [Lab15], has become a powerful alternative for scheduling. This is evidenced by a CP model for resource-constrained project scheduling provided by [LRSV18]. Additionally, a recent comparison between solvers for CP and MIP demonstrated CP Optimizer’s superiority over CPLEX across a broad set of benchmark scheduling problems [NRR23]. These advances, however, have not yet been explored for SRCPSP/max, despite potential applications. CP can be used for finding robust proactive schedules or within a reactive approach with rescheduling during execution, which has been viewed as too computationally heavy [VdVDH07].
A proper benchmarking paper in which a statistical analysis is performed on the results of the different methods for SRCPSP/max is lacking. Comparing different stochastic scheduling techniques for this problem should be done carefully. Due to the maximal time lags and stochastic durations, methods can fail by violating resource or precedence constraints. Therefore, not only the solution quality and computation time but also the feasibility should be taken into account. [FVL16] even criticize their own evaluation metric and indicate that future work should provide a more extensive comparison. We argue that statistical tests dealing with infeasibilities correctly are needed to compare different methods properly.
This paper proposes new versions of a proactive approach and a hybrid approach using the latest advances in Constraint Programming [LRSV18] and Temporal Networks [HP24]. A new reactive approach with complete rescheduling, which has not been investigated before as far as we know, is included in the comparison too. We target the existing research gap due to the lack of a clear and informative comparison among various methods, examining aspects such as infeasibility, solution quality, and computation time (offline and online). In doing so, this work provides a practical guide for researchers and industrial schedulers when selecting from various methods for their specific use cases.
2 The Scheduling Problem
The RCPSP/max problem [KS96] is defined as a set of activities $J$ and a set of resources $R$, where each activity $j\in J$ has a duration $d_{j}$ and requires from each resource $r\in R$ a certain amount indicated by $r_{r,j}$. A solution to the RCPSP/max is a start-time assignment $s_{j}$ for each activity, such that the following constraints are satisfied: (i) at any time the number of resources used cannot exceed the max resource capacity $c_{r}$; and (ii) for some pairs $(i,j)$ of activities, precedence constraints are defined as minimal or maximal time lags between the start time of $i$ and the start time of $j$. The goal is to minimize the makespan of the schedule.
We use the following example from [SFSW13] of a simple RCPSP/max instance. There are five activities: $a$, $b$, $c$, $d$, and $e$, with start times $s_{a}$, $s_{b}$, $s_{c}$, $s_{d}$, and $s_{e}$. The activities have durations of 2, 5, 3, 1, and 2 time units, respectively, and resource requirements of 3, 2, 1, 2, and 2 on a single resource with a capacity of $c_{r}=4$. The precedence relations are as follows: activity $b$ starts at least 2 time units after $a$ starts, activity $b$ starts at least 1 time unit before $c$ starts, activity $c$ cannot start later than 6 time units after $a$ starts, and activity $d$ starts exactly 3 time units before $e$ starts. These constraints can be visualized in a project graph (Figure 1). A feasible solution to this problem is $\{s_{a}=1,s_{b}=3,s_{c}=4,s_{d}=0,s_{e}=3\}$, for which the Gannt chart is visualized in Figure 2 (ignore the arrows for now).
Figure 1: Example of project graph, figure from [SFSW13].
A special property of (S)RCPSP/max is the following:
Proposition 1.
Suppose we are given a problem instance $1$ with durations $d^{1}$, resource requirements $r^{1}$, and capacity $c^{1}$, and let $s^{1}$ be a feasible schedule for this instance. Suppose now that we transform instance $1$ into instance $2$, where all parameters stay equal except that one or more of the activity durations $d^{2}$ are shorter than the durations $d^{1}$, so $\forall\ j\in J:d^{1}_{j}\leq d^{2}_{j}$. Then, $s^{1}$ is also feasible for instance $2$.
Proof.
Schedule $s^{1}$ is still precedence feasible because the start times did not change and the precedence constraints are defined from start to start (they are deterministic). Schedule $s^{1}$ is resource feasible for $d^{1}$. Since $d^{2}$ is strictly smaller than $d^{1}$, the resource usage over time can only be smaller than for instance $1$, and thus it will not exceed the capacity, and schedule $s^{1}$ is also resource feasible for $d^{2}$. ∎
This property can be helpful when reusing start time solutions for a mutated instance, for example, because of stochastic activity durations. SRCPSP/max is an extension in which the durations follow a stochastic distribution [FLVX12]. The durations $d_{j}$ are random variables bounded by an uncertainty interval and will become available when an activity finishes. Earlier work on this stochastic variant introduces the $\alpha$-robust makespan [FLVX12], that is the expected makespan value for a scheduling strategy for which the probability that the schedule is feasible is at least $1-\alpha$, as a metric for comparing different methods.
3 Background and Related Work
In this section, we discuss existing approaches for SRCPSP/max. Furthermore, we introduce all concepts that are needed to understand our scheduling methods that are presented in Section 4. We introduce proactive techniques based on Sample Average Approximation in Section 3.1. We explain partial order schedules and temporal networks in Section 3.2. Finally, Section 3.3 gives an overview of related work on benchmarking scheduling methods for SRCPSP/max.
3.1 Proactive Scheduling
Proactive scheduling methods aim to find a robust schedule offline by taking information about the uncertainty into account [HL02]. In this section, we explain a core technique used in proactive methods: Sample Average Approximation (SAA). Furthermore, we discuss the state-of-the-art proactive methods for SRCPSP/max.
3.1.1 Sample Average Approximation
A common method for handling discrete optimization under uncertainty is the Sample Average Approximation (SAA) approach [KSHdM02]. Samples are drawn from stochastic distributions and added as scenarios to a stochastic programming formulation. The solver then seeks a solution feasible for all scenarios while optimizing the average objective. However, adding more samples to the SAA increases the number of constraints and variables, significantly raising the solution time.
3.1.2 SORU and SORU-H
The most recent proactive method on SRCPSP/max is proposed by [VFL16] and is recognized as the state of the art. The authors present the algorithm $\mathsf{SORU}$, an SAA approach to the scheduling problem that relies on Mixed Integer Programming (MIP) and aims to minimize the $\alpha$-robust makespan. It can be summarized as follows: (i) a selection of samples is used to set up the SAA; (ii) the model seeks a start time vector $s$ such that the minimal and maximal time lags of precedence constraints are satisfied; (iii) it allows for $\alpha\%$ of the scenarios to be resource infeasible; (iv) it minimizes the sample average makespan.
Since $\mathsf{SORU}$ is computationally expensive, the authors also propose a heuristic version, dubbed $\mathsf{SORU}$-$\mathsf{H}$. Instead of a set of samples, one single summarizing sample is used, which represents a quantile of the distribution. Note that because of Proposition 1, this heuristic approximates the $\alpha$-robust makespan. At the same time, it is much cheaper to compute because typically the runtime increases for a larger sample size in SAA, as shown by [VFL16].
Since nowadays CP is the state of the art for deterministic project scheduling [NRR23], we re-investigate SAA approaches for SRCPSP/max with CP. We expect that the scalability of SAA is much less of a problem than it was when the $\mathsf{SORU}$ methods were presented.
3.2 Partial Order Scheduling
In this section, we discuss partial order scheduling approaches, which can be seen as a reactive-proactive hybrid. Partial order schedules have been used in the majority of the contributions to SRCPSP/max.
3.2.1 Constructing Ordering Constraints
A partial order schedule (POS) can be seen as a collection of schedules that ensure resource feasibility, but maintain temporal flexibility. A POS is defined as a graph where nodes represent activities, and edges temporal constraints between them. There are several methods to derive a POS. In the original paper by [PSCO04], two approaches are outlined to construct the ordering constraints between activities, either analyzing the resource profile to avoid all possible resource conflicts, based on Minimal Critical Sets (MCS) [LR83], or using a fixed resource-feasible schedule together with a chaining procedure to construct resource chains (see Figure 2). They find that using the single-point solution together with chaining seems both simple and most effective. Subsequent work explored alternative MCS-based approaches [LM09, LMB13] and chaining heuristics [FLVX12].
Figure 2: Example Gannt chart figure adjusted from [FLVX12]. The X-axis is time, and Y-axis shows resource demand. The arrows indicate an example of generating a POS with chaining, starting from a fixed solution schedule.
3.2.2 Temporal Networks
Partial order schedules are often complemented with a temporal model to reason over the temporal constraints. Most POS approaches rely on a Simple Temporal Network (STN), which is a graph consisting of time points (nodes) and temporal difference constraints (edges). The Simple Temporal Network with Uncertainty (STNU) extends the STN by introducing contingent links. The duration of these contingent constraints can only be observed, while for regular constraints it can be determined during execution. The works by [LM09] and [LMB13] on POS are the only ones that we know of using STNUs as their temporal model. They introduce nodes for the start and the end of each activity with duration constraints between them, connecting respective start nodes with edges for the precedence constraints.
3.2.3 Execution Strategies
An STNU is dynamically controllable (DC) if there is a strategy to determine execution times for all controllable (non-contingent) time points which ensures that all temporal constraints are met, regardless of the outcomes of the contingent links. Some DC-checking algorithms generate new so-called wait edges to make the network DC [Mor14]. The DC STNU with wait edges is referred to as an Extended STNU (ESTNU) and can be given to a Real-Time Execution Algorithm. Such an algorithm is the online component that transforms an STNU into a schedule (i.e. an execution time for each time point), given the observations for the contingent time points. An algorithm specifically tailored to ESTNUs is RTE [HP24].
As far as we know, the DC-checking and RTE algorithms have not been applied to SRCPSP/max, despite their efficiency. [LM09, LMB13] used constraint propagation for DC-checking, but do not use RTE. Other works on POS for SRCPSP/max that do not use STNUs and DC-checking risk violations of minimal or maximal time-lags during execution [PCOS07, FVL16]. Thus, we conclude that there is a research gap in applying the developments in STNU literature to SRCPSP/max.
3.3 Benchmarking Approaches
Benchmarking procedures for comparing scheduling methods are inconsistent in the literature, with varying problem sets and comparison methods. Some studies used industrial scheduling instances [LM09, LMB13], while the majority [PSCO04, FLVX12, VFL16, FVL16] relied on PSPlib [KS96] instances, which are deterministic and transformed into stochastic versions with noise. Research typically focused on instances with 10, 20, and 30 activities (j10-j30). Different studies assessed varying metrics, such as schedule flexibility and robustness [PSCO04, PCOS07], solver performance [LM09, LMB13], or the $\alpha$-robust makespan [FLVX12, VFL16, FVL16]. However, no comprehensive benchmarking paper exists that evaluates both solution quality and computation time while also correctly accounting for infeasibilities. We take inspiration from [LF03], who compare different planners that can fail, providing a framework for comparing solution quality and speed while also correctly considering correctly considering these failures.
4 Scheduling Methods
This section outlines the proposed methods. We explain how to use CP for these scheduling problems so far dominated by MIP approaches.
First, we present a deterministic CP model for RCPSP/max in Section 4.1. The new, stochastic methods for SRCPSP/max are proposed in Section 4.2. We explain the statistical tests for performing pairwise comparisons of these new methods that lead to partial orderings based on solution quality and computation time in Section 4.3.
4.1 Constraint Programming for RCPSP/max
The CP model is:
[…]
$\sum_{j\in J}\mathrm{Pulse}(x_{j},r_{r,j})\leq c_{r}$ $\ \forall r\in R$ (1e)
$x_{j}:\mathrm{IntervalVar}(J,d_{j})$ $\ \forall j\in J$ (1f)
For the deterministic RCPSP/max, we use the modern interval constraints from the IBM CP optimizer [te17]. The CP model is defined in equations 1a-1f , we use IBM’s syntax and modify the RCPSP example from [LRSV18] to RCPSP/max. We use the earlier introduced nomenclature together with the minimal time lags $min_{j,i}$ and maximal time lags $max_{j,i}$ that are the temporal differences between start times of activities $j$ and $i$ if $i$ is a successor of $j$. We introduce the decision variable $x_{j}$ as the interval variable for activity $j\in J$. The $\mathrm{Pulse}$ function generates a cumulative expression over a given interval $x_{j}$ with a certain value. For a task $j$, this value is its resource usage $r_{r,j}$. The aggregated pulse values are constrained so that their sum does not exceed the total available resource capacity $c_{r}$.
4.2 New Methods for SRCPSP/max
This section introduces three new methods for SRCPSP/max. The first method is a CP-based version of a proactive model. Then, we present a novel, fully reactive scheduling approach employing the deterministic model for RCPSP/max. Finally, we propose an STNU-based approach using CP and POS. We refer to these approaches as $\mathsf{proactive}$, $\mathsf{reactive}$, and $\mathsf{stnu}$, respectively.
4.2.1 Proactive Method
We outline how to use a scenario-based CP model (instead of MIP) for SRCPSP/max which we call $\mathsf{proactive}_{\mathsf{SAA}}$.
For this SAA method, we can reuse the deterministic RCPSP/max CP model, introduced in Section 4.1, but we introduce scenarios. This model is inspired by the MIP version by [VFL16]. A special variant is the SAA with only one sample, for which a $\gamma$-quantile can be used which we call $\mathsf{proactive}_{\gamma}$. If a feasible schedule can be found for the $\gamma$-quantile, this schedule will also be feasible for all duration realizations on the left-hand side of the $\gamma$-quantile because of Proposition 1. We provide the SAA model below. We use the same nomenclature as for the deterministic model, but we introduce the notion of scenarios $\omega\in\Omega$, and find a schedule $s_{j}\ \forall j\in J$ that is feasible for all scenarios it has seen if one exists.
[…]
$\sum_{j\in J}\mathrm{Pulse(x_{j}^{\omega},r_{r,j})\leq c_{r}}$ $\ \forall r\in R\ \forall\omega\in\Omega$ (2e)
$x_{j}^{\omega}:\mathrm{IntervalVar(J,y_{j}^{\omega})}$ $\ \forall j\in J\ \forall\omega\in\Omega$ (2f)
$s_{j}==\mathrm{start}(x_{j}^{\omega})$ $\ \forall j\in J\ \forall\omega\in\Omega$ (2g)
4.2.2 Reactive Method
We now present a CP-based fully reactive approach for SRCPSP/max that we refer to as $\mathsf{reactive}$.
Fully reactive approaches, involving complete rescheduling by solving a deterministic RCPSP/max, have been considered impractical due to high computational demands and low schedule stability [VdVDH07]. However, advances in CP for scheduling [LRSV18, NRR23] mitigate these issues. In practice, decision-makers often reschedule their entire future plan when changes occur. Thus, including this reactive approach in our comparison is valuable. To our knowledge, such an approach has not been evaluated before. The outline is:
  • Start by making an initial schedule with an estimation of the activity durations $\hat{d}$. We can see $\hat{d}$ as a hyperparameter for how conservative the estimation is. For example, using the mean of the distribution could lead to better makespans, but the risk of becoming infeasible for larger duration realizations, while taking the upper bound of the distribution could lead to a very high makespan.
  • At every decision moment (i.e. when an activity finishes) resolve the deterministic RCPSP/max while fixing all variables until the current time to reschedule with new information, we again use the estimation $\hat{d}$ for the activities that did not finish yet. Resolving is only needed when the finish time of an activity deviates from the estimated finish time. We warm start the solver with the previous solution.
4.2.3 STNU-based Method
Finally, we present a partial order schedule approach using CP and STNU algorithms [HP24] which we call $\mathsf{stnu}$. This approach is inspired by many earlier works [PCOS07, LM09, FVL16]. We use a fixed-solution approach for constructing the ordering constraints. The outline of the approach is:
  1. 1.
    We make a fixed-point schedule by solving the deterministic RCPSP/max with an estimation of the activity durations $\widehat{d}$ and using the chaining procedure, which was explained in Section 3.2.
  2. 2.
    We construct the STNU as follows:
    • For each activity two nodes are created, representing its start and end.
[…]
The resulting STNU is tested for dynamic controllability (DC), and if the network is DC its extended form (ESTNU) is given to the RTE algorithm. Since makespan minimization is of interest, we slightly adjust the algorithm from [HP24]: instead of selecting an arbitrary executable time point and an arbitrary allowed execution time, we always choose the earliest possible time point at the earliest possible execution time.
4.3 Statistical Tests for Pairwise Comparison
A strategy for benchmarking is to provide partial orderings of the scheduling methods for the different metrics of interest (i.e. solution quality, runtime offline, runtime online). A partial ordering can be obtained by executing pairwise comparisons of the methods per problem size and per metric, taking inspiration from [LF03].
The Wilcoxon Matched-Pairs Rank-Sum Test (the version by [Cur67]) looks at the ranking of absolute differences and gives insight into which of a pair of methods has consistently better performance than another method. Infeasible cases can be handled by assigning infinitely bad time and solution quality to these cases, leading to an absolute difference of $\infty$ or $\infty$- ∞ that will be pushed to the highest and lowest rankings.
An alternative test to the Wilcoxon test that is also used by [LF03] is the proportion test (see Test 4 in the book by [Kan06]). This test is weaker than the Wilcoxon, but provides at least information about significance in the proportion of wins when the Wilcoxon test shows no significant difference.
An interesting additional test is to test whether there is a significant difference in the magnitude of the different metrics. This can be tested with a pairwise t-test on two related samples of scores (see Test 10 in the book by [Kan06]). This test can only be performed on double hits, because infinitely bad computation time or solution quality will disturb the test.
We provide a detailed explanation of all of the above statistical tests in our Technical Appendix.
Test Legend $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
Wilc. Quality [n] z (p) [370] -6.75 (*) [370] -6.8 (*) [370] -6.86 (*) [370] -2.76 (*) [370] -8.53 (*) [370] -7.09 (*)
Prop. Quality [n] prop (p) [277] 0.78 (*) [276] 0.78 (*) [278] 0.78(*) [61] 0.67 (*) [73] 1.0 (*) [65] 0.94 (*)
Magn. Quality [n] t (p) [330] -11.36 (*) [330] -11.39 (*) [330] -11.35 (*) [370] -2.86 (*) [370] -6.73 (*) [370] -6.01 (*)
norm. avg. $\mathsf{stnu}$: 0.985 $\mathsf{stnu}$: 0.985 $\mathsf{stnu}$: 0.984 $\mathsf{react}$: 1.0 $\mathsf{react}$: 0.999 $\mathsf{pro}_{\mathsf{SAA}}$: 0.999
norm. avg. $\mathsf{react}$: 1.015 $\mathsf{pro}_{\mathsf{SAA}}$: 1.015 $\mathsf{pro}_{0.9}$: 1.016 $\mathsf{pro}_{\mathsf{SAA}}$: 1.0 $\mathsf{pro}_{0.9}$: 1.001 $\mathsf{pro}_{0.9}$: 1.001
Test Legend $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$
Wilc. Offline [n] z (p) [370] -16.67 (*) [370] -16.67 (*) [370] -16.67 (*) [370] -16.67 (*) [370] -7.26 (*)
Prop. Offline [n] prop (p) [370] 1.0 (*) [370] 1.0 (*) [370] 1.0 (*) [370] 1.0 (*) [370] 0.62 (*)
Magn. Offline [n] t (p) [330] -36.15 (*) [370] -72.56 (*) [330] -36.15 (*) [370] -72.56 (*) [330] -5.64 (*)
norm. avg. $\mathsf{react}$: 0.36 $\mathsf{react}$: 0.24 $\mathsf{pro}_{0.9}$: 0.36 $\mathsf{pro}_{0.9}$: 0.24 $\mathsf{stnu}$: 0.82
norm. avg. $\mathsf{stnu}$: 1.64 $\mathsf{pro}_{\mathsf{SAA}}$: 1.76 $\mathsf{stnu}$: 1.64 $\mathsf{pro}_{\mathsf{SAA}}$: 1.76 $\mathsf{pro}_{\mathsf{SAA}}$: 1.18
Test Legend $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
Wilc. Online [n] z (p) [370] -16.67 (*) [370] -16.67 (*) [370] -16.67 (*) [370] -16.67 (*) [370] -9.85 (*)
Prop. Online [n] prop (p) [370] 1.0 (*) [370] 1.0 (*) [370] 1.0 (*) [370] 1.0 (*) [370] 0.89 (*)
[…]
norm. avg. $\mathsf{pro}_{0.9}$: 0.01 $\mathsf{pro}_{\mathsf{SAA}}$: 0.01 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.15
norm. avg. $\mathsf{stnu}$: 1.99 $\mathsf{stnu}$: 1.99 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.85
Table 1: Pairwise test results for the instance set ubo50, noise level $c=1$, covering solution quality, offline and online runtime. For each metric, we show the test results of the Wilcoxon test, proportion test, and magnitude test such as described in Section 4.3. Wilcoxon test results are presented as: [nr pairs] z-value (p-value), with (*) for $p<0.05$. Proportion test results are: [nr pairs] proportion (p-value). Magnitude test results are: [nr pairs] t-stat (p-value), with normalized averages of the two methods in the last two rows. Additional results for $c\in[1,2]$, j10, j20, j30, ubo50, and ub0100 are in the Technical Appendix. Each column header lists the two methods compared, with the first being the better performer, excluding pairs with only zero-differences.
5 Experimental Evaluation
The goal of the evaluation is to analyze the relative performance of the different scheduling methods proposed in our methodology.
5.1 Data Generation
We use the j10, j20, j30, ubo50, and ubo100 sets from the PSPlib [KS96]. Instead of using all instances from j10 to j30, we select 50 per set and extend previous work [VFL16, FVL16] by including 50 instances each from the ubo50 and ubo100 sets. Following prior research, we use deterministic durations to set up stochastic distributions with different noise levels, converting deterministic instances into stochastic ones. We use uniform discrete distributions for durations based on the deterministic processing times $d_{j}$. Specifically, we define the lower bound as $lb=\max(1,\text{round}(d_{j}-c\cdot\sqrt{d_{j}}))$ and the upper bound as $ub=\text{round}(d_{j}+c\cdot\sqrt{d_{j}})$, where we vary the noise level $c=\{1,2\}$. The source and sink nodes always have a deterministic duration of zero. Each evaluation of a method corresponds to one sample from the distribution, and we sample $10$ times for each instance. We excluded the instance samples for which it is impossible to find a feasible schedule (i.e. the deterministic problem with perfect information is infeasible for the given activity duration sample).
5.2 Tuning of the Methods
In this section, we highlight the most important observations and outcomes of our tuning results; for further details, see our Technical Appendix.
All CP models are solved with the IBM CP solver [te17] with default settings. Most of the deterministic RCPSP/max the j10-j30 instances can be solved within 60 seconds. For, ubo50 and ubo100 instances, we fixed the time limit to 600 seconds. We tune the classical SAA with multiple scenarios $\mathsf{proactive}_{\mathsf{SAA}}$ and a heuristic approach $\mathsf{proactive}_{\gamma}$. We used $\gamma=0.9$ for $\mathsf{proactive}_{\gamma}$. For $\mathsf{proactive}_{\mathsf{SAA}}$, we sampled quantiles at $\gamma\in[0.25,0.5,0.75,0.9]$ and set a time limit of 1800 seconds. The algorithm $\mathsf{reactive}$ uses offline the $\mathsf{proactive}_{\gamma}$ algorithm, for which we fixed $\gamma=0.9$. We investigated the effect of the time limit for rescheduling and set this hyperparameter to 2 seconds. The algorithm $\mathsf{stnu}$ consists of an offline procedure, which comprises building up the network and checking dynamic controllability; if it’s not DC, the method fails. We observed that that the choice of the $\gamma$-quantile for the fixed-point schedule significantly affects the ratio of DC networks. We select $\gamma=1$ for our final method $\mathsf{stnu}$, because this setting results in much more DC networks than lower quantiles.
5.3 Results
We include an analysis of the feasibility ratios of the different methods. A selection of the results of the statistical tests are shown in Table 1 (the remaining results are included in the Technical Appendix). We present summarized partial orderings on 1) solution quality, 2) time offline, 3) time online.
5.3.1 Feasibility Ratio
We first analyze the ratio of solved problems per problem set and method in Table 10. For $c=1$, the feasibility ratios of $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$, and $\mathsf{reactive}$ are similar for instance sets j10-ubo50, with $\mathsf{reactive}$ slightly lower for ubo100. The $\mathsf{stnu}$ method has the lowest feasibility rate, but the difference narrows as problem size increases. For $c=2$, the $\mathsf{stnu}$ achieves the highest feasibility ratios due to better handling of larger variances in durations.
c, Set $\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$ $\mathsf{reactive}$
1, j10 0.65 0.85 0.85 0.85
1, j20 0.65 0.76 0.76 0.76
1, j30 0.78 0.89 0.89 0.89
1, ubo50 0.77 0.86 0.86 0.86
1, ubo100 0.84 0.91 0.91 0.88
2, j10 0.63 0.63 0.64 0.63
2, j20 0.63 0.54 0.53 0.53
2, j30 0.63 0.51 0.48 0.49
2, ubo50 0.67 0.41 0.42 0.41
2, ubo100 0.79 0.35 0.36 0.33
Table 2: Feasibility ratio.
5.3.2 Results Statistical Tests
Table 1 shows a subset of the pairwise test results for the metrics solution quality, runtime offline, and runtime online. In our Technical Appendix, we provide all test results per instance set / noise level $c$ setting. The outcomes of the test results can be used to make partial orderings of the methods, distinguishing between a strong partial ordering (Wilcoxon test) and a weak ordering (proportion test).
5.3.3 Partial Ordering Visualization
In Figure 3 - 5, an arrow $A\rightarrow B$ indicates that in the majority of the settings either $A$ is consistently better than $B$ (Wilcoxon) and/or $A$ is better than $B$ a significant number of times (proportion). Due to space constraints and for clarity, we have chosen to display only the most common pattern per metric rather than all partial orderings per instance set and noise level. Consequently, the distinction between strong and weak partial orderings is omitted in these figures. We refer to the Technical Appendix for the partial orderings per setting including the distinction between a strong and weak partial ordering.
5.4 Analysis
Figure 3 shows the visualization of the partial ordering for solution quality (makespan). The results are consistent for the different instance sets and noise levels. The $\mathsf{stnu}$ shows to be the outperforming method based on solution quality. The $\mathsf{reactive}$ approach outperforms the $\mathsf{proactive}$ methods. Furthermore, $\mathsf{proactive}_{\mathsf{SAA}}$ outperforms in many cases $\mathsf{proactive}_{0.9}$, although for larger instances and a higher noise level a significant difference is not present. In earlier work, $\mathsf{proactive}$ approaches were considered state-of-the-art, but in our analysis, we found better makespan results for the STNU-based approach. We found that for each pair for which an arrow is visualized in Figure 3 also a significant magnitude difference was found on the double hits.
Figure 3: Summarizing illustration of the partial ordering of the different methods for solution quality.
Figure 4 shows that the $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$ have the lowest relative offline runtime and we found no significant difference between the two. The ordering of $\mathsf{stnu}$ and $\mathsf{proactive}_{\mathsf{SAA}}$ depends on the problem size (the ordering flips for larger instances). These relative orderings are also confirmed with the magnitude test on double hits. However, we assign infinitely bad offline computation time to infeasible solutions. When executing the Wilcoxon test we include all instances for which at least one of the two methods generated a feasible solution. For that reason, a flip in the partial ordering occurs for $c=2$, ubo50 and ubo100: $\mathsf{stnu}$ shows the best performance, and $\mathsf{proactive}_{0.9}$ outperforms $\mathsf{reactive}$ according to the Wilcoxon tests. This was mainly due to the higher feasibility ratio for the better methods (see Table 10) as the results from the magnitude test on double hits contradicted Wilcoxon in these cases (we did not visualize this pattern in the main paper, but we included it in the Appendix).
Figure 4: Summarizing illustration of the partial ordering of the different methods for time offline.
We observe the general partial ordering for online runtime in Figure 5. The superiority of $\mathsf{proactive}_{0.9}$ and $\mathsf{proactive}_{\mathsf{SAA}}$ are expected, as these methods only require a feasibility check online. The faster online time of the $\mathsf{stnu}$ compared to the $\mathsf{reactive}$ can be explained by the fact that the $\mathsf{stnu}$ employs a polynomial real-time execution algorithm, while $\mathsf{reactive}$ calls a deterministic CP solver multiple times. These results are confirmed with a magnitude test on double hits. Again, there are a few settings ($c=2$, ubo50 and ubo100) in which the magnitude and the Wilcoxon test contradict each other: $\mathsf{stnu}$ outperforms the $\mathsf{proactive}$ methods based on the Wilcoxon test due to higher feasibility ratios, while the magnitude test on double hits shows better online runtime for $\mathsf{proactive}_{0.9}$ and $\mathsf{proactive}_{\mathsf{SAA}}$.
Figure 5: Summarizing illustration of the partial ordering of the different methods for time offline.
5.5 Reproducibility
Together with this article, we provide a Technical Appendix in which we include explanations of the statistical tests, hyperparameter tuning, tables with test results from all pairwise comparisons that led to the partial orderings in Figure 3, 4 and 5. Furthermore, we provide a zip file comprising our code that can be used to rerun all experiments and statistical tests. In the camera-ready version of this paper, this section will refer to a public GitHub repository of our code.
6 Conclusion and Future Work
This study proposes new scheduling methods for SRCPSP/max using the latest advances in CP and STNUs. We conducted an extensive, statistical analysis to compare the proposed algorithms for the problem instances and target the existing research gap of a lacking benchmarking paper for this problem.
Until now, a MIP-based proactive method was considered the best method for SRCPSP/max, although partial order schedules have shown potential in earlier research. We found that proactive methods can be improved with online rescheduling, resulting in better solution quality for the method $\mathsf{reactive}$ compared to proactive approaches $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{proactive}_{\gamma}$. We find that the algorithm $\mathsf{stnu}$ that uses partial order schedules outperforms the other methods on solution quality in our evaluation. Although in general, $\mathsf{proactive}_{\gamma}$ and $\mathsf{reactive}$ have better offline computation time than $\mathsf{stnu}$, and $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{proactive}_{\gamma}$ have better online computation time than $\mathsf{stnu}$, the $\mathsf{stnu}$ also showed good relative runtime results due to the polynomial time STNU-related algorithms.
In future work, the same approach could be used to evaluate other scheduling problems, and we can gain more insight into how these methods perform on both well-known problems from the literature and in practical situations. Furthermore, the set of methods could even be broadened by including sequential methods such as those provided by [Hal22] and/or machine learning-based approaches such as the graph neural network approach by [TKPdGB+23] that was developed for stochastic RCPSP.
Appendix A Tuning
This section describes the tuning process for the different scheduling methods. First, we investigated the effect of the time limit on solving the deterministic CP. Then, for the proactive approach, we tuned the time limit and the sample selection for the Sample Average Approximation (SAA). For $\mathsf{reactive}$, we tuned the quantile approximation and the time limit for rescheduling.
A.1 Time limit for CP on deterministic instances
To understand how the time limit may affect the results, we first consider the deterministic instances of RCPSP/max. [VFL16] reported on their time budget that $\mathsf{SORU}$ was able to obtain solutions within 5 minutes for every one instance in J10 and 2 hrs for the J20 instances. However, for J30, we were unable to get optimal solutions for certain instances in the cut-off limit of 3 hrs. On the other hand, $\mathsf{SORU}$-$\mathsf{H}$ was able to generate solutions for J10 instances within half of a second, J20 instances within 10 seconds, and J30 instances within 10 minutes on average.”
In our research, we extend the instance sets with ubo50 and ubo100, for which the time limit can become more crucial. For each set (j10-ubo100), we fixed the first 50 instances from the PSPlib [KS96] (PSP1 - PSP50). For j10-j30, the IBM CP Optimizer could solve all instances (PSP1-PSP50) within 60 seconds to optimality or prove infeasibility.
For the ubo50 instances PSP1 to PSP10 and the ubo100 instances PSP1 to PSP10, the effect of the time limits of 60 seconds, 600 seconds, and 3600 seconds respectively, is presented in Table 3 and 4. We observe that the solver status does not change when increasing the time limit to 600 seconds, and only flips 1 instance from feasible to infeasible when increasing the time limit from 600 seconds to 3600 seconds. The makespan only improves significantly for the ubo100 instances for higher time limits. In the remaining experiments, we fixed the time limit to 60 seconds for solving deterministic CPs.
Instance set Time limit Feasible Infeasible Optimal Unkown
ubo50 60s 3 3 2 2
ubo50 600s 3 3 2 2
ubo50 3600s 2 3 3 2
ubo100 60s 4 5 0 1
ubo100 600s 4 5 0 1
ubo100 3600s 4 5 0 1
Table 3: Count per solver status IBM CP solver after different time limits on deterministic instances.
60s 600s 3600s
ubo50 204 200 200
ubo100 421 410 410
Table 4: Average makespan (filtered feasible and optimal solutions on ubo50 and ubo100 set.
A.2 Proactive Approach
The sample selection process is expected to influence the performance of the proactive method. We used a subset of the j10 and ubo50 instances (for both sets $PSP\_1$-$PSP\_20$, and used 10 duration samples per instance. We fixed the time limit to $60$ seconds and compared the feasibility ratios in Table 5. We found that the higher feasibility ratios were obtained using two settings for the proactive method, being 1 sample with $\gamma=0.9$, to which we will refer with $\mathsf{proactive}_{0.9}$, and the setting with the smart samples to which we will refer with $\mathsf{proactive}_{\mathsf{SAA}}$ in the remaining of our experiments. For the SAA with four samples (smart samples), we investigated the effect of the time limit on the makespan in Table 6. We observed that the makespan would still improve while making the step from 600 seconds to 3600 seconds. We decide to use a time limit of 1800 seconds instead of 3600 seconds in the remaining of the experiments to be able to conduct more experiments.
1 sample $\gamma=$0.5 $\gamma=$0.75 $\gamma=$0.9 $\gamma=$1
j10 0.11 0.53 0.94 0.75
ubo50 0 0.01 0.92 0.80
multiple samples n=5 n=10 n=25 n=50
j10 0.63 0.78 0.94 0.94
ubo50 0.12 0.48 0.89 0.91
smart samples* n=4
j10 0.94
ubo50 0.92
Table 5: Feasibility ratio for different $\gamma$-quantiles in proactive approach. Each cell counts hits out of 200 experiments. * Uses a combination of smart samples that are the quantiles with $\gamma\in[0.25,0.5,0.75,0.9]$
60s 600s 3600s
ubo50 233 232 231
ubo100 485 480 472
Table 6: Average makespan obtained with $\mathsf{proactive}_{\mathsf{SAA}}$ with four scenarios for different time limits (filtered feasible and optimal solutions on ubo50 and ubo100 set.
A.3 Reactive Approach
First, we observed the effect of the duration estimations on the performance of $\mathsf{reactive}$. We used a subset of the j10 and ubo50 instances (for both sets $PSP\_1$-$PSP\_20$, and used 10 duration samples per instance. We fixed the time limit for the initial schedule to $60$ seconds and for the rescheduling to $2$ seconds.
$\gamma=$0.5 $\gamma=$0.75 $\gamma=$0.9 $\gamma=$1
j10 0.11 0.53 0.94 0.75
ubo50 0 0.01 0.92 0.80
Table 7: Feasibility ratio for different $\gamma$-quantiles in reactive approach. Each cell counts hits out of 200 experiments.
Remarkably, we observe similar feasibility ratios to the proactive approach, indicating that for feasibility the initial schedule is quite important. For the final evaluation, we fixed $\gamma=0.9$ for $\mathsf{reactive}$, and will analyze how the solution quality improves with the rescheduling procedure compared to a standard proactive approach.
Next, we observe the effect of the time limit for rescheduling. We used the same subset of the j10 and ubo50 instances (for both sets $PSP\_1$-$PSP\_20$, used 10 duration samples per instance, and runtime limits of 1, 2, 10 and 30 seconds. The results (in Table 7) are almost similar for the different time limits, this might be because the solver finishes already before the time limit, and the increase in time online has mainly to do with the number of solver calls. In the experimental evaluation, we therefore fixed the rescheduling time limit to $2$ seconds, which we expected to be sufficient for larger or slightly more complicated problems.
time limit rescheduling 1s 2s 10s 30s
j10 makespan 38 38 38 39
j10 time online 0.03 0.04 0.04 0.04
ubo50 makespan 171 171 171 172
ubo 50 time online 15.4 15.1 15.1 15.4
Table 8: Average makespan and time online for double hits on different time limits for rescheduling. Each cell averages over double hits out of 200 experiments.
A.4 Hyperparameters Selection
This subsection presents the hyperparameters in Table 9 that are used in the final experiments that are also presented in the main paper.
Hyperparameter $\mathsf{proactive}_{\mathsf{SAA}}$ $\mathsf{proactive}_{0.9}$ $\mathsf{stnu}$ $\mathsf{reactive}$
$\gamma$ [0.25, 0.5, 0.75, 0.7] 0.9 1 0.9
Time limit CP 1800s 600s 600s 600s and 2s
Solver IBM CP IBM CP IBM CP IBM CP
Table 9: Hyperparameter settings.
Appendix B Statistical Tests
B.1 Wilcoxon Test
The Wilcoxon test that we use follows [Cur67] and is described below:
  • Collect a set of matched pairs (the results from two different methods on one instance sample).
  • Compute the difference between the two test results for each pair.
    An important remark is that because of the discrete objective values (makespan), we can obtain a zero difference when there is a tie, we use Pratt’s procedure for handling ties [Pra59], which includes zero-differences in the ranking process, but drops the ranks of the zeros afterward.
  • Order the pairs according to the absolute values of the differences.
  • Assign ranks to the pairs based on these absolute values.
  • Sum the positive ranks ($T_{pos})$ and the negative ($T_{neg})$ ranks separately.
  • Take the smaller of the two $T=min(T_{pos},T_{neg})$.
  • If the two methods have no consistent difference in their relative performances, then the rank-sums should be approximately equal. This is tested with a normal approximation for the Wilcoxon statistic which is outlined by [Cur67]. [Cur67] propose a corrected normal approximation which is needed because of usage of the the Pratt procedure for handling zero differences.
  • The normalized Z-statistic is given by the formula: $z=(T-d)/se$, where $d$ is the continuity correction from [Cur67], and $se$ is the standard error.
All of the above can be executed using the Python package SciPy [VGO+20] built-in method $scipy.stats.wilcoxon$ that is called with parameters $method$=”approx”, ”$zero\_method$=”pratt”, and $correction$=True.
B.2 Z-test for Proportion (Binomial Distribution)
We use [Kan06] as a reference for the Z-test for Proportion. It is important to mention that this test is approximate as it assumes that the number of observations justifies a normal approximation for the binomial. (In contradiction to the SciPy package and its built-in method $scipy.stats.binomtest$ containing an exact test).
The proportion test investigates whether the is a significant difference between the assumed proportion of wins $p_{0}$ and the observed proportion of wins $p$. In our analysis, two methods are compared and the number of wins for each method is counted based on one metric.
The procedure for the proportion test is as follows:
  • Collect a set of matched pairs (the results from two different methods on one atomic instance form a pair).
  • For each pair, determine which method wins, and count the wins for both methods. Exclude all ties.
  • Calculate the ratio of wins for one of the two methods.
  • Test where this ratio differs significantly from $p_{0}=0.5$ (equal probability of winning).
    • The test statistic is $Z=\frac{|p-p_{0}|-\frac{1}{2n}}{\{\frac{p_{0}(1-p_{0})}{n}^{1/2}\}}$
    • The term $\frac{1}{2n}$ in the numerator is a discontinuity correction.
    • For a two-sided test with a significance level $\alpha=0.05$ the acceptance region for the null hypothesis is $-1.96<Z<1.96$.
B.3 Magnitude Test
The magnitude test we use is a t-test for two population means, or the method of paired comparisons such as Test 10 in the book by [Kan06]. The test whether there is a significant difference between two population means. The procedure for this paired comparison t-test is as follows:
  • Collect a set of matched pairs (so the results from two different methods on one atomic instance forms a pair).
  • Normalise the performances for each pair by computing the mean value of the pair and dividing the two items in the pair by the pairs’ mean such that all normalized observations will be between 0 and 2, and 1 indicates a tie.
  • Compute the differences $d_{1}$ between the two test results for each pair $i$.
  • Compute the variances of the differences with the following formula: $s^{2}=\sum_{i=1}^{n}\frac{(d_{i}-\bar{d})^{2}}{n-1}.$
  • Compute the means of both methods $\bar{x_{1}}$ and $\bar{x_{2}}$.
  • Compute the t-statistic using the formula: $t=\frac{\bar{x_{1}}-\bar{x_{2}}}{s/n^{1/2}}$
  • Test significance by checking whether $t$ lies within the acceptance region for which the values are given by the Student’s t-distribution (two-sided) with $n-1$ degrees of freedom.
After the normalization step, it is possible to execute the test with the Scipy package and specifically, its built-in method $scipy.stats.ttest\_rel$.
Appendix C Results Tables
Please find the results of the statistical test in this section. This data led to the Figures that are included in our main paper.
  • Tables 12 for noise level $c=1$ and 14 for noise level $c=2$: including the results of the Wilcoxon and proportion tests on solution quality.
  • Tables 13 for noise level $c=1$ and 15 for noise level $c=2$: including the results of the magnitude test (independent t-test) on solution quality.
  • Tables 16 for noise level $c=1$ and 18 for noise level $c=2$: including the results of the Wilcoxon and proportion tests on time offline.
  • Tables 17 for noise level $c=1$ and 19 for noise level $c=2$: including the results of the magnitude test (independent t-test) on time offline.
  • Tables 20 and Tables 22: including the results of the Wilcoxon and proportion tests on time online.
  • Tables 21 and 23: including the results of the magnitude test (independent t-test) on time online.
C.1 Feasibility Ratio
First, we analyze the feasibility ratio obtained by the different methods. For $c=1$, we observe the feasibility ratio obtained by $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$, $\mathsf{reactive}$ are the same for instance sets j10 - ubo50, and only for ubo100 $\mathsf{reactive}$ is a little bit lower than the $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$. The $\mathsf{stnu}$ method has the lowest feasibility rate but remarkably the difference becomes smaller when the size of the problem grows (the difference is the smallest for ubo100).
We observe a different pattern for $c=2$, with a higher noise factor. The highest feasibility ratios are obtained by $\mathsf{stnu}$. This could be explained by the fact that the $\mathsf{stnu}$ method uses the information about the distribution and is, therefore, better at handling the larger variances in activity duration.
Set $\mathsf{stnu}$ $\mathsf{proactive}_{\mathsf{SAA}}$ $\mathsf{proactive}_{0.9}$ $\mathsf{reactive}$
j10 0.65 0.85 0.85 0.85
j20 0.65 0.76 0.76 0.76
j30 0.78 0.89 0.89 0.89
ubo50 0.77 0.86 0.86 0.86
ubo100 0.84 0.91 0.91 0.88
Table 10: Feasibility ratio for noise factor $c=1$ We use abbreviations for the proactive approach with $\gamma=0.9$ ($\mathsf{pro}_{0.9}$), and the SAA variant ($\mathsf{pro}_{\mathsf{SAA}}$).
Set $\mathsf{stnu}$ $\mathsf{proactive}_{\mathsf{SAA}}$ $\mathsf{proactive}_{0.9}$ $\mathsf{reactive}$
j10 0.63 0.63 0.64 0.63
j20 0.63 0.54 0.53 0.53
j30 0.63 0.51 0.48 0.49
ubo50 0.67 0.41 0.42 0.41
ubo100 0.79 0.35 0.36 0.33
Table 11: Feasibility ratio for noise factor $c=2$ We use abbreviations for the proactive approach with $\gamma=0.9$ ($\mathsf{pro}_{0.9}$), and the SAA variant ($\mathsf{pro}_{\mathsf{SAA}}$).
C.2 Partial Ordering Solution Quality
Next, we analyze the test results based on the Wilcoxon test, proportion test, and magnitude test for solution quality. We translate the test results into partial orderings that are visualized with solid arrows indicating a significant difference based on the Wilcoxon test and dashed arrows indicating a significant difference based on the proportion test for the pairs where the Wilcoxon test did not find significant results. Furthermore, for all pairs where we find a significant difference according to Wilcoxon, or the proportion test, we analyze whether we can find a significant magnitude difference on the double hits.
  • First, we discuss partial orderings for the noise level $c=1$, which are visualized in Figure 6. For $c=1$, the partial ordering shows the same pattern across the different instance sets. However, there are a few exceptions:
    • In some cases we can only show a dashed arrow, indicating a weaker partial ordering, that is a significant proportional difference (proportion of wins) instead of a significant difference obtained by Wilcoxon.
    • Only for ubo50, all arrows are solid, indicating that for all pairs we found a significant difference according to the Wilcoxon test.
    • For ubo100, there is no significant difference between $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{proactive}_{0.9}$. Possibly for the larger instances the SAA becomes more difficult to solve, for which reason it is not better than the heuristic $\gamma$-quantile procedure anymore.
  • Then, we analyze the partial orderings for the noise level $c=2$, in Figure 7. Here there are a few things to observe:
    • We observe a shift in the partial orderings for $c=2$.
    • For j10-j20, the partial ordering is similar to $c=1$, although for $c=2$ all arrows are solid (indicating a relation obtained by the Wilcoxon test).
    • For $j30$, there is no significant difference between $\mathsf{proactive}_{0.9}$ and $\mathsf{proactive}_{\mathsf{SAA}}$, neither with Wilcoxon and neither proportionally.
    • The partial orderings for $j30$ and $ubo50$ are similar.
    • For ubo100, we observe the $\mathsf{stnu}$ is significantly better than the three other methods, although no significant difference can be found between the three other methods.
  • The main conclusion is that in all situations the $\mathsf{stnu}$ shows to be the outperforming method based on solution quality. In general, the $\mathsf{reactive}$ shows to outperform the $\mathsf{proactive}$methods. Furthermore, $\mathsf{proactive}_{\mathsf{SAA}}$ outperforms in many cases $\mathsf{proactive}_{0.9}$, although for larger instances and a higher noise level this difference is not present anymore.
Figure 6: Partial ordering on solution quality for $c=1$. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
Figure 7: Partial ordering on solution quality for $c=2$. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
C.3 Partial Ordering Time Offline
  • We expected the $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$ to be the fastest offline, followed by the $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{stnu}$, where the partial ordering between the latter two depends on the problem size. When two methods have (almost) only ties (such as $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$) not all tests can be executed and a $nan$ will appear.
  • Importantly, the tests are designed in such a way that infeasibilities are also weighted in significance testing. When a method results in an infeasible solution, infinite offline time is assigned to this experiment. For that reason, it can still occur that we find a difference between $\mathsf{reactive}$ and $\mathsf{proactive}_{0.9}$: although they have the same offline procedure, one of the two can still fail which results in infinite offline time.
  • In general, we give preference to the test results of the Wilcoxon test because this test shows a stronger relation than the proportional test, but to fill in the missing information we use the proportion test to test if there is a significant difference in the proportion of wins. We noted, however, that the test outcomes of the two tests can sometimes be contradicting. The Wilcoxon test penalizes infeasibilities more severely, while the proportion test only considers the number of wins. Consequently, it is possible that method 1 produces more infeasible solutions but achieves better metrics, whereas method 2 generates fewer infeasible solutions but worse metrics, resulting in the Wilcoxon test on instances where at least one feasible solution exists tells us method 2 performs better (infeasible solutions are penalized more heavily). Proportion test on instances where at least one feasible solution exists tells us method 1 performs better. The magnitude test on double hits tells us Method 1 performs better as it has better metric values.
  • Now, we observe the different partial orderings obtained for $c=1$:
    • In general, the time spent offline for the $\mathsf{proactive}_{0.9}$ and the $\mathsf{reactive}$ is exactly the same, therefore in most cases we also did not find a significant difference between the two. However, especially for the ubo100 instances, we found that sometimes $\mathsf{proactive}_{0.9}$ had more feasible solutions, resulting in a partial ordering of $\mathsf{proactive}_{0.9}\rightarrow\mathsf{reactive}$.
    • We observe that for j10 and j20 $\mathsf{proactive}_{\mathsf{SAA}}$ is consistently faster offline than $\mathsf{stnu}$, but this flips for j30 - ubo100, where the $\mathsf{stnu}$ becomes faster. This can be explained by the fact that the solve time of the $SAA$ can grow exponentially, while the DC-checking algorithm is a polynomial time algorithm.
  • For noise level $c=2$, we observe in Figure 9:
    • For j10 and j20, the partial ordering is almost the same as for the $c=1$ level, although the flipping already occurs at j20, where the $\mathsf{stnu}$ becomes faster than the $\mathsf{proactive}_{\mathsf{SAA}}$.
    • Surprisingly, the pattern changes for the ubo50 and ubo100 instances. We observe that the best performance on time offline is obtained by the $\mathsf{stnu}$ according to the Wilcoxon. We found that this is mainly caused by the much higher feasibility ratio of the $\mathsf{stnu}$ method.
    • Furthermore, we find that the $\mathsf{proactive}_{0.9}$ becomes better than $\mathsf{reactive}$ because of the higher feasibility ratio again (as the two methods have the same offline procedure).
    • Again, for the larger instances sets (j20-ubo100), the $\mathsf{proactive}_{\mathsf{SAA}}$ has the worst relative computation time offline.
Figure 8: Partial ordering time offline c=1. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
Figure 9: Partial ordering time offline c=2. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
C.4 Partial Ordering Time Online
We observe the results from the tests on online computation time:
  • For $c=1$ (Figure 11), we find the same partial ordering for each instance set, which is also in line with our expectations:
    • There is no difference between $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{proactive}_{0.9}$ as both only require a feasibility check only.
    • The $\mathsf{stnu}$ real-time execution algorithm turns out to be more efficient online than the reactive method, which comprises multiple solver calls, this is also expected.
  • For $c=2$ (Figure 11), we find the same pattern as for $c=1$, for problem sets j10-j30, however, we find that for ubo50 and ubo100, the $\mathsf{stnu}$ starts to outperform other methods, explained by the higher feasibility ratio of the $\mathsf{stnu}$.
Figure 10: Partial ordering time online c=1, j10, j20, j30, ubo50, and ubo100. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
Figure 11: Partial ordering time online c=2, ubo50, and ubo100. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
C.5 Magnitude Tests
In general, we observed that for the pairs of methods for which we found a significant partial ordering, we also found a significant difference in the magnitude of the different metrics. The test results for solution quality are presented in Table 13 (c=1) and 15 (c=2), for time offline in Table 17 (c=1) and 19 (c=2).
The remarkable things that we observed in the results of the magnitude tests, were also already observed while comparing Wilcoxon test results with proportional test results in the earlier analysis for a few cases:
  • For time offline, $c=2$, ubo50 and ubo100, the Wilcoxon test found that the $\mathsf{stnu}$ is better than e.g. $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$, while looking at the magnitude results on double hits (this is an important difference in test procedure!), we find that the $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$ are much faster offline than $\mathsf{stnu}$, which is also expected. Then, we observe no difference between $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$, which can be explained by the fact that both have the same offline procedure and we select only double hits for the test.
  • We find something similar occurs for the time online with $c=2$, according to Wilcoxon, $\mathsf{stnu}$ is outperforming $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{proactive}_{0.9}$ for ubo50 and ubo100. However, this is again mainly caused by the higher ratio of feasible solutions, as when we analyze the magnitude difference the proactive methods are much faster online looking at the double hits (which is logical as this comprises only of feasibility checking).
C.6 Summary
In our main paper, we employed the statistical analysis from this document to provide a summarizing partial ordering per metric including a brief summary of the main findings. We decide to present the most occuring patterns, and in text highlight any important exceptions.
Appendix D Reproducibility
We uploaded our code to make it possible to reproduce the results of our research and to facilitate others using a similar benchmarking method. Please refer to the READ.ME of our repository to find instructions for requirements and directions to our experiment scripts, and the scripts to run the statistical tests.
j10 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[340] -1.659 (0.097) [340] -0.067 (0.947) [340] -0.082 (0.935) [340] -10.571 (*) [340] -10.918 (*) [340] -5.756 (*)
[308] 0.591 (*) [310] 0.632 (*) [310] 0.632 (*) [117] 0.991 (*) [121] 1.0 (*) [37] 0.973 (*)
j20 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[270] -1.523 (0.128) [270] -2.401 (*) [270] -2.679 (*) [270] -5.936 (*) [270] -7.852 (*) [270] -6.144 (*)
[213] 0.596 (*) [215] 0.633 (*) [213] 0.643 (*) [75] 0.827 (*) [62] 1.0 (*) [59] 0.898 (*)
j30 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[320] -0.892 (0.372) [320] -1.218 (0.223) [320] -2.157 (*) [320] -5.54 (*) [320] -9.395 (*) [320] -5.299 (*)
[214] 0.579 (*) [214] 0.575 (*) [217] 0.618 (*) [103] 0.757 (*) [89] 1.0 (*) [79] 0.797 (*)
ubo50 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[370] -6.748 (*) [370] -6.8 (*) [370] -6.859 (*) [370] -2.76 (*) [370] -8.531 (*) [370] -7.085 (*)
[277] 0.776 (*) [276] 0.779 (*) [278] 0.781 (*) [61] 0.672 (*) [73] 1.0 (*) [65] 0.938 (*)
ubo100 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[400] -2.12 (*) [400] -1.76 (0.078) [400] -1.842 (0.065) [400] -0.442 (0.659) [400] -6.534 (*) [400] -0.2 (0.842)
[312] 0.519 (0.533) [314] 0.51 (0.778) [311] 0.514 (0.65) [147] 0.51 (0.869) [88] 0.864 (*) [151] 0.503 (1.0)
Table 12: Pairwise comparison on schedule quality (makespan) for noise factor c=1. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each paier method 1 - method 2. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] z-value (p-value) with (*) for p $<$ 0.05
j10 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[260] -9.367 (*) [260] -13.762 (*) [260] -13.703 (*) [340] -10.114 (*) [340] -10.542 (*) [340] -5.637 (*)
$\mathsf{stnu}$: 0.974 $\mathsf{stnu}$: 0.962 $\mathsf{stnu}$: 0.962 $\mathsf{react}$: 0.986 $\mathsf{react}$: 0.985 $\mathsf{pro}_{\mathsf{SAA}}$: 0.999
$\mathsf{react}$: 1.026 $\mathsf{pro}_{\mathsf{SAA}}$: 1.038 $\mathsf{pro}_{0.9}$: 1.038 $\mathsf{pro}_{\mathsf{SAA}}$: 1.014 $\mathsf{pro}_{0.9}$: 1.015 $\mathsf{pro}_{0.9}$: 1.001
j20 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[230] -7.62 (*) [230] -8.608 (*) [230] -8.992 (*) [270] -6.868 (*) [270] -7.713 (*) [270] -5.82 (*)
$\mathsf{stnu}$: 0.982 $\mathsf{stnu}$: 0.978 $\mathsf{stnu}$: 0.977 $\mathsf{react}$: 0.995 $\mathsf{react}$: 0.993 $\mathsf{pro}_{\mathsf{SAA}}$: 0.998
$\mathsf{react}$: 1.018 $\mathsf{pro}_{\mathsf{SAA}}$: 1.022 $\mathsf{pro}_{0.9}$: 1.023 $\mathsf{pro}_{\mathsf{SAA}}$: 1.005 $\mathsf{pro}_{0.9}$: 1.007 $\mathsf{pro}_{0.9}$: 1.002
j30 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[280] -4.072 (*) [280] -5.946 (*) [280] -7.06 (*) [320] -5.698 (*) [320] -9.422 (*) [320] -6.226 (*)
$\mathsf{stnu}$: 0.993 $\mathsf{stnu}$: 0.99 $\mathsf{stnu}$: 0.988 $\mathsf{react}$: 0.997 $\mathsf{react}$: 0.995 $\mathsf{pro}_{\mathsf{SAA}}$: 0.998
$\mathsf{react}$: 1.007 $\mathsf{pro}_{\mathsf{SAA}}$: 1.01 $\mathsf{pro}_{0.9}$: 1.012 $\mathsf{pro}_{\mathsf{SAA}}$: 1.003 $\mathsf{pro}_{0.9}$: 1.005 $\mathsf{pro}_{0.9}$: 1.002
ubo50 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[330] -11.359 (*) [330] -11.391 (*) [330] -11.346 (*) [370] -2.858 (*) [370] -6.734 (*) [370] -6.014 (*)
$\mathsf{stnu}$: 0.985 $\mathsf{stnu}$: 0.985 $\mathsf{stnu}$: 0.984 $\mathsf{react}$: 1.0 $\mathsf{react}$: 0.999 $\mathsf{pro}_{\mathsf{SAA}}$: 0.999
$\mathsf{react}$: 1.015 $\mathsf{pro}_{\mathsf{SAA}}$: 1.015 $\mathsf{pro}_{0.9}$: 1.016 $\mathsf{pro}_{\mathsf{SAA}}$: 1.0 $\mathsf{pro}_{0.9}$: 1.001 $\mathsf{pro}_{0.9}$: 1.001
ubo100 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[358] -6.316 (*) [370] -6.813 (*) [370] -6.817 (*) [388] -2.833 (*) [388] -7.738 (*) [400] -0.514 (0.608)
$\mathsf{stnu}$: 0.991 $\mathsf{stnu}$: 0.99 $\mathsf{stnu}$: 0.99 $\mathsf{react}$: 0.999 $\mathsf{react}$: 0.999 $\mathsf{pro}_{\mathsf{SAA}}$: 1.0
$\mathsf{react}$: 1.009 $\mathsf{pro}_{\mathsf{SAA}}$: 1.01 $\mathsf{pro}_{0.9}$: 1.01 $\mathsf{pro}_{\mathsf{SAA}}$: 1.001 $\mathsf{pro}_{0.9}$: 1.001 $\mathsf{pro}_{0.9}$: 1.0
Table 13: Magnitude test on solution quality for noise factor c=1. Using a pairwise t-test, including all instances for which both methods found a feasible solution. Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.
j10 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[258] -5.194 (*) [258] -7.825 (*) [258] -7.822 (*) [235] -10.082 (*) [234] -11.046 (*) [235] -4.058 (*)
[237] 0.696 (*) [247] 0.802 (*) [248] 0.802 (*) [146] 0.911 (*) [134] 0.993 (*) [37] 0.838 (*)
j20 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[217] -9.825 (*) [217] -10.194 (*) [216] -10.57 (*) [182] -4.58 (*) [177] -7.243 (*) [182] -3.241 (*)
[185] 0.854 (*) [188] 0.872 (*) [188] 0.888 (*) [79] 0.759 (*) [53] 1.0 (*) [60] 0.7 (*)
j30 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[233] -6.986 (*) [234] -7.175 (*) [232] -7.993 (*) [183] -1.355 (0.175) [163] -7.557 (*) [182] -5.689 (*)
[191] 0.749 (*) [192] 0.776 (*) [196] 0.796 (*) [94] 0.574 (0.18) [58] 1.0 (*) [96] 0.802 (*)
ubo50 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[284] -11.76 (*) [285] -11.73 (*) [284] -11.799 (*) [171] -1.973 (*) [162] -3.74 (*) [173] -0.938 (0.348)
[236] 0.907 (*) [238] 0.908 (*) [236] 0.911 (*) [50] 0.66 (*) [39] 0.821 (*) [36] 0.583 (0.405)
ubo100 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[285] -12.139 (*) [284] -11.55 (*) [288] -11.26 (*) [141] -1.83 (0.067) [122] -1.017 (0.309) [146] -0.083 (0.934)
[261] 0.851 (*) [254] 0.799 (*) [265] 0.796 (*) [82] 0.415 (0.151) [44] 0.477 (0.88) [82] 0.5 (0.912)
Table 14: Pairwise comparison on schedule quality (makespan) for noise factor c=2. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] z-value (p-value) with (*) for p $<$ 0.05 for the proportion test.
j10 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[205] -9.128 (*) [204] -16.872 (*) [206] -17.395 (*) [230] -10.858 (*) [233] -11.811 (*) [231] -4.856 (*)
$\mathsf{stnu}$: 0.963 $\mathsf{stnu}$: 0.929 $\mathsf{stnu}$: 0.925 $\mathsf{react}$: 0.966 $\mathsf{react}$: 0.964 $\mathsf{pro}_{\mathsf{SAA}}$: 0.997
$\mathsf{react}$: 1.037 $\mathsf{pro}_{\mathsf{SAA}}$: 1.071 $\mathsf{pro}_{0.9}$: 1.075 $\mathsf{pro}_{\mathsf{SAA}}$: 1.034 $\mathsf{pro}_{0.9}$: 1.036 $\mathsf{pro}_{0.9}$: 1.003
j20 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[170] -10.855 (*) [172] -11.237 (*) [170] -11.89 (*) [174] -4.95 (*) [176] -6.22 (*) [173] -3.656 (*)
$\mathsf{stnu}$: 0.972 $\mathsf{stnu}$: 0.968 $\mathsf{stnu}$: 0.964 $\mathsf{react}$: 0.991 $\mathsf{react}$: 0.989 $\mathsf{pro}_{\mathsf{SAA}}$: 0.997
$\mathsf{react}$: 1.028 $\mathsf{pro}_{\mathsf{SAA}}$: 1.032 $\mathsf{pro}_{0.9}$: 1.036 $\mathsf{pro}_{\mathsf{SAA}}$: 1.009 $\mathsf{pro}_{0.9}$: 1.011 $\mathsf{pro}_{0.9}$: 1.003
j30 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[140] -4.151 (*) [148] -6.017 (*) [138] -6.149 (*) [152] -4.199 (*) [160] -6.746 (*) [150] -5.404 (*)
$\mathsf{stnu}$: 0.99 $\mathsf{stnu}$: 0.986 $\mathsf{stnu}$: 0.985 $\mathsf{react}$: 0.994 $\mathsf{react}$: 0.991 $\mathsf{pro}_{\mathsf{SAA}}$: 0.996
$\mathsf{react}$: 1.01 $\mathsf{pro}_{\mathsf{SAA}}$: 1.014 $\mathsf{pro}_{0.9}$: 1.015 $\mathsf{pro}_{\mathsf{SAA}}$: 1.006 $\mathsf{pro}_{0.9}$: 1.009 $\mathsf{pro}_{0.9}$: 1.004
ubo50 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[141] -7.974 (*) [146] -8.139 (*) [148] -8.125 (*) [145] -3.581 (*) [155] -4.646 (*) [150] -2.823 (*)
$\mathsf{stnu}$: 0.984 $\mathsf{stnu}$: 0.981 $\mathsf{stnu}$: 0.979 $\mathsf{react}$: 0.998 $\mathsf{react}$: 0.997 $\mathsf{pro}_{\mathsf{SAA}}$: 0.999
$\mathsf{react}$: 1.016 $\mathsf{pro}_{\mathsf{SAA}}$: 1.019 $\mathsf{pro}_{0.9}$: 1.021 $\mathsf{pro}_{\mathsf{SAA}}$: 1.002 $\mathsf{pro}_{0.9}$: 1.003 $\mathsf{pro}_{0.9}$: 1.001
ubo100 $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$
[94] -0.305 (0.761) [114] -0.296 (0.767) [114] -0.93 (0.355) [76] -1.573 (0.12) [99] -4.187 (*) [94] 0.59 (0.556)
$\mathsf{stnu}$: 0.999 $\mathsf{stnu}$: 0.999 $\mathsf{stnu}$: 0.998 $\mathsf{react}$: 0.999 $\mathsf{react}$: 0.999 $\mathsf{pro}_{\mathsf{SAA}}$: 1.0
$\mathsf{react}$: 1.001 $\mathsf{pro}_{\mathsf{SAA}}$: 1.001 $\mathsf{pro}_{0.9}$: 1.002 $\mathsf{pro}_{\mathsf{SAA}}$: 1.001 $\mathsf{pro}_{0.9}$: 1.001 $\mathsf{pro}_{0.9}$: 1.0
Table 15: Magnitude test on solution quality for noise factor c=2. Using a pairwise t-test, including all instances for which both methods found a feasible solution. Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.
j10 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$
[340] nan (nan) [340] -15.982 (*) [340] -15.982 (*) [340] -15.982 (*) [340] -15.982 (*) [340] -13.986 (*)
[nan] nan (nan) [340] 1.0 (*) [340] 1.0 (*) [340] 1.0 (*) [340] 1.0 (*) [340] 0.882 (*)
j20 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$
[270] nan (nan) [270] -14.125 (*) [270] -14.245 (*) [270] -14.125 (*) [270] -14.245 (*) [270] -3.927 (*)
[nan] nan (nan) [270] 0.963 (*) [270] 1.0 (*) [270] 0.963 (*) [270] 1.0 (*) [270] 0.63 (*)
j30 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$
[320] nan (nan) [320] -13.963 (*) [320] -15.506 (*) [320] -13.963 (*) [320] -15.506 (*) [320] -1.63 (0.103)
[nan] nan (nan) [320] 0.969 (*) [320] 1.0 (*) [320] 0.969 (*) [320] 1.0 (*) [320] 0.5 (0.955)
ubo50 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$
[370] nan (nan) [370] -16.671 (*) [370] -16.671 (*) [370] -16.671 (*) [370] -16.671 (*) [370] -7.261 (*)
[nan] nan (nan) [370] 1.0 (*) [370] 1.0 (*) [370] 1.0 (*) [370] 1.0 (*) [370] 0.622 (*)
ubo100 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$
[400] -3.463 (*) [400] -14.988 (*) [400] -15.286 (*) [400] -15.969 (*) [400] -17.332 (*) [400] -6.406 (*)
[12] 1.0 (*) [400] 0.962 (*) [400] 0.97 (*) [400] 0.975 (*) [400] 1.0 (*) [400] 0.6 (*)
Table 16: Pairwise comparison on time offline for noise factor c=1. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2 according to Wilcoxon. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] ratio of wins (p-value) with (*) for p $<$ 0.05
j10 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$
[340] nan (nan) [260] -365.854 (*) [340] -77.22 (*) [260] -365.854 (*) [340] -77.22 (*) [260] -19.399 (*)
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.14 $\mathsf{react}$: 0.32 $\mathsf{pro}_{0.9}$: 0.14 $\mathsf{pro}_{0.9}$: 0.32 $\mathsf{pro}_{\mathsf{SAA}}$: 0.65
$\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.86 $\mathsf{pro}_{\mathsf{SAA}}$: 1.68 $\mathsf{stnu}$: 1.86 $\mathsf{pro}_{\mathsf{SAA}}$: 1.68 $\mathsf{stnu}$: 1.35
j20 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$
[270] nan (nan) [230] -45.769 (*) [270] -52.707 (*) [230] -45.769 (*) [270] -52.707 (*) [230] -3.062 (*)
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.22 $\mathsf{react}$: 0.28 $\mathsf{pro}_{0.9}$: 0.22 $\mathsf{pro}_{0.9}$: 0.28 $\mathsf{pro}_{\mathsf{SAA}}$: 0.91
$\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.78 $\mathsf{pro}_{\mathsf{SAA}}$: 1.72 $\mathsf{stnu}$: 1.78 $\mathsf{pro}_{\mathsf{SAA}}$: 1.72 $\mathsf{stnu}$: 1.09
j30 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ stnu-$\mathsf{pro}_{\mathsf{SAA}}$
[320] nan (nan) [280] -38.703 (*) [320] -53.836 (*) [280] -38.703 (*) [320] -53.836 (*) [280] 0.5 (0.618)
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.3 $\mathsf{react}$: 0.32 $\mathsf{pro}_{0.9}$: 0.3 $\mathsf{pro}_{0.9}$: 0.32 $\mathsf{stnu}$: 1.02
$\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.7 $\mathsf{pro}_{\mathsf{SAA}}$: 1.68 $\mathsf{stnu}$: 1.7 $\mathsf{pro}_{\mathsf{SAA}}$: 1.68 $\mathsf{pro}_{\mathsf{SAA}}$: 0.98
ubo50 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ stnu-$\mathsf{pro}_{\mathsf{SAA}}$
[370] nan (nan) [330] -36.15 (*) [370] -72.563 (*) [330] -36.15 (*) [370] -72.563 (*) [330] -5.636 (*)
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.36 $\mathsf{react}$: 0.24 $\mathsf{pro}_{0.9}$: 0.36 $\mathsf{pro}_{0.9}$: 0.24 $\mathsf{stnu}$: 0.82
$\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.64 $\mathsf{pro}_{\mathsf{SAA}}$: 1.76 $\mathsf{stnu}$: 1.64 $\mathsf{pro}_{\mathsf{SAA}}$: 1.76 $\mathsf{pro}_{\mathsf{SAA}}$: 1.18
ubo100 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$
[388] nan (nan) [358] -27.209 (*) [388] -70.349 (*) [370] -25.293 (*) [400] -70.926 (*) [370] -6.391 (*)
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.43 $\mathsf{react}$: 0.27 $\mathsf{pro}_{0.9}$: 0.45 $\mathsf{pro}_{0.9}$: 0.27 $\mathsf{stnu}$: 0.79
$\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.57 $\mathsf{pro}_{\mathsf{SAA}}$: 1.73 $\mathsf{stnu}$: 1.55 $\mathsf{pro}_{\mathsf{SAA}}$: 1.73 $\mathsf{pro}_{\mathsf{SAA}}$: 1.21
Table 17: Magnitude test on time offline for noise factor c=1. Using a pairwise t-test, including all instances for which both methods found a feasible solution. Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.
j10 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$
[234] -0.996 (0.319) [258] -9.382 (*) [235] -12.848 (*) [258] -9.554 (*) [235] -13.069 (*) [258] -6.664 (*)
[nan] nan (nan) [258] 0.903 (*) [235] 0.991 (*) [258] 0.907 (*) [235] 0.996 (*) [258] 0.764 (*)
j20 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$
[177] -0.994 (0.32) [217] -4.543 (*) [182] -10.458 (*) [216] -4.456 (*) [182] -10.213 (*) [217] -5.281 (*)
[nan] nan (nan) [217] 0.816 (*) [182] 0.973 (*) [216] 0.815 (*) [182] 0.967 (*) [217] 0.604 (*)
j30 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$
[163] -1.728 (0.084) [233] -0.083 (0.934) [183] -7.137 (*) [232] -0.267 (0.789) [182] -6.667 (*) [234] -7.974 (*)
[3] 0.0 (0.248) [233] 0.7 (*) [183] 0.891 (*) [232] 0.69 (*) [182] 0.879 (*) [234] 0.726 (*)
ubo50 $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$
[284] -4.827 (*) [284] -5.557 (*) [285] -10.588 (*) [162] -2.643 (*) [173] -8.741 (*) [171] -7.484 (*)
[284] 0.454 (0.138) [284] 0.479 (0.514) [285] 0.712 (*) [7] 1.0 (*) [173] 0.936 (*) [171] 0.906 (*)
ubo100 $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$
[288] -8.783 (*) [285] -10.58 (*) [284] -13.068 (*) [122] -4.786 (*) [146] -5.008 (*) [141] -1.441 (0.15)
[288] 0.576 (*) [285] 0.653 (*) [284] 0.82 (*) [23] 1.0 (*) [146] 0.836 (*) [141] 0.702 (*)
Table 18: Pairwise comparison on time offline for noise factor c=2. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2 according to Wilcoxon. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] ratio of wins (p-value) with (*) for p $<$ 0.05
j10 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$
[233] nan (nan) [205] -229.689 (*) [230] -59.39 (*) [206] -230.812 (*) [231] -59.404 (*) [204] -17.066 (*)
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.15 $\mathsf{react}$: 0.35 $\mathsf{pro}_{0.9}$: 0.15 $\mathsf{pro}_{0.9}$: 0.35 $\mathsf{pro}_{\mathsf{SAA}}$: 0.61
$\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.85 $\mathsf{pro}_{\mathsf{SAA}}$: 1.65 $\mathsf{stnu}$: 1.85 $\mathsf{pro}_{\mathsf{SAA}}$: 1.65 $\mathsf{stnu}$: 1.39
j20 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$
[176] nan (nan) [170] -82.607 (*) [174] -63.732 (*) [170] -82.607 (*) [173] -63.382 (*) [172] 0.379 (0.705)
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.21 $\mathsf{react}$: 0.23 $\mathsf{pro}_{0.9}$: 0.21 $\mathsf{pro}_{0.9}$: 0.23 $\mathsf{stnu}$: 1.01
$\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.79 $\mathsf{pro}_{\mathsf{SAA}}$: 1.77 $\mathsf{stnu}$: 1.79 $\mathsf{pro}_{\mathsf{SAA}}$: 1.77 $\mathsf{pro}_{\mathsf{SAA}}$: 0.99
j30 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$
[160] nan (nan) [140] -33.765 (*) [152] -40.336 (*) [138] -33.253 (*) [150] -39.904 (*) [148] -2.168 (*)
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.34 $\mathsf{react}$: 0.33 $\mathsf{pro}_{0.9}$: 0.34 $\mathsf{pro}_{0.9}$: 0.33 $\mathsf{stnu}$: 0.9
$\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.66 $\mathsf{pro}_{\mathsf{SAA}}$: 1.67 $\mathsf{stnu}$: 1.66 $\mathsf{pro}_{\mathsf{SAA}}$: 1.67 $\mathsf{pro}_{\mathsf{SAA}}$: 1.1
ubo50 $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$
[148] 26.653 (*) [141] 27.168 (*) [146] -0.952 (0.343) [155] nan (nan) [150] -45.457 (*) [145] -44.29 (*)
$\mathsf{stnu}$: 1.7 $\mathsf{stnu}$: 1.72 $\mathsf{stnu}$: 0.95 $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{pro}_{0.9}$: 0.24 $\mathsf{react}$: 0.24
$\mathsf{pro}_{0.9}$: 0.3 $\mathsf{react}$: 0.28 $\mathsf{pro}_{\mathsf{SAA}}$: 1.05 $\mathsf{react}$: 1.0 $\mathsf{pro}_{\mathsf{SAA}}$: 1.76 $\mathsf{pro}_{\mathsf{SAA}}$: 1.76
ubo100 $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{stnu}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$
[114] 14.765 (*) [94] 15.183 (*) [114] -3.125 (*) [99] nan (nan) [94] -42.151 (*) [76] -42.944 (*)
$\mathsf{stnu}$: 1.55 $\mathsf{stnu}$: 1.61 $\mathsf{stnu}$: 0.83 $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{pro}_{0.9}$: 0.21 $\mathsf{react}$: 0.18
$\mathsf{pro}_{0.9}$: 0.45 $\mathsf{react}$: 0.39 $\mathsf{pro}_{\mathsf{SAA}}$: 1.17 $\mathsf{react}$: 1.0 $\mathsf{pro}_{\mathsf{SAA}}$: 1.79 $\mathsf{pro}_{\mathsf{SAA}}$: 1.82
Table 19: Magnitude test on time offline for noise factor c=2. Using a pairwise t-test, including all instances for which both methods found a feasible solution. Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.
j10 $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[340] -0.368 (0.713) [340] -15.993 (*) [340] -15.997 (*) [340] -15.98 (*) [340] -15.98 (*) [340] -2.726 (*)
[85] 0.471 (0.664) [340] 1.0 (*) [340] 1.0 (*) [340] 1.0 (*) [340] 1.0 (*) [340] 0.765 (*)
j20 $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[270] -1.17 (0.242) [270] -14.246 (*) [270] -14.245 (*) [270] -14.243 (*) [270] -14.243 (*) [270] -6.441 (*)
[133] 0.556 (0.225) [270] 1.0 (*) [270] 1.0 (*) [270] 1.0 (*) [270] 1.0 (*) [270] 0.852 (*)
j30 $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[320] -0.653 (0.514) [320] -15.505 (*) [320] -15.505 (*) [320] -15.504 (*) [320] -15.504 (*) [320] -8.247 (*)
[191] 0.524 (0.563) [320] 1.0 (*) [320] 1.0 (*) [320] 1.0 (*) [320] 1.0 (*) [320] 0.875 (*)
ubo50 $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[370] -0.771 (0.44) [370] -16.67 (*) [370] -16.67 (*) [370] -16.67 (*) [370] -16.67 (*) [370] -9.859 (*)
[299] 0.522 (0.488) [370] 1.0 (*) [370] 1.0 (*) [370] 1.0 (*) [370] 1.0 (*) [370] 0.892 (*)
ubo100 $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[400] -0.719 (0.472) [400] -17.331 (*) [400] -17.331 (*) [400] -17.331 (*) [400] -17.331 (*) [400] -12.488 (*)
[391] 0.453 (0.069) [400] 1.0 (*) [400] 1.0 (*) [400] 1.0 (*) [400] 1.0 (*) [400] 0.925 (*)
Table 20: Pairwise comparison on time online for noise factor c=1. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] proportion (p-value) with (*) for p $<$ 0.05
j10 $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[340] 0.679 (0.498) [260] -109.014 (*) [260] -106.917 (*) [340] -4393.475 (*) [340] -4277.995 (*) [260] -1151.565 (*)
$\mathsf{pro}_{0.9}$: 1.02 $\mathsf{pro}_{0.9}$: 0.05 $\mathsf{pro}_{\mathsf{SAA}}$: 0.05 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.04
$\mathsf{pro}_{\mathsf{SAA}}$: 0.98 $\mathsf{stnu}$: 1.95 $\mathsf{stnu}$: 1.95 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.96
j20 $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[270] -0.663 (0.508) [230] -285.563 (*) [230] -282.464 (*) [270] -7311.952 (*) [270] -7470.984 (*) [230] -605.38 (*)
$\mathsf{pro}_{0.9}$: 0.97 $\mathsf{pro}_{0.9}$: 0.03 $\mathsf{pro}_{\mathsf{SAA}}$: 0.03 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.07
$\mathsf{pro}_{\mathsf{SAA}}$: 1.03 $\mathsf{stnu}$: 1.97 $\mathsf{stnu}$: 1.97 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.93
j30 $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[320] -0.156 (0.876) [280] -789.246 (*) [280] -786.726 (*) [320] -14638.229 (*) [320] -14425.009 (*) [280] -338.664 (*)
$\mathsf{pro}_{0.9}$: 0.99 $\mathsf{pro}_{0.9}$: 0.02 $\mathsf{pro}_{\mathsf{SAA}}$: 0.02 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.09
$\mathsf{pro}_{\mathsf{SAA}}$: 1.01 $\mathsf{stnu}$: 1.98 $\mathsf{stnu}$: 1.98 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.91
ubo50 $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[370] -0.795 (0.427) [330] -3951.584 (*) [330] -3976.441 (*) [370] -30723.506 (*) [370] -31447.766 (*) [330] -142.758 (*)
$\mathsf{pro}_{\mathsf{SAA}}$: 0.97 $\mathsf{pro}_{0.9}$: 0.01 $\mathsf{pro}_{\mathsf{SAA}}$: 0.01 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.15
$\mathsf{pro}_{0.9}$: 1.03 $\mathsf{stnu}$: 1.99 $\mathsf{stnu}$: 1.99 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.85
ubo100 $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[400] -0.065 (0.948) [370] -56517.571 (*) [370] -48810.689 (*) [388] -83435.86 (*) [388] -78013.468 (*) [358] -60.202 (*)
$\mathsf{pro}_{\mathsf{SAA}}$: 1.0 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.23
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{stnu}$: 2.0 $\mathsf{stnu}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.77
Table 21: Magnitude t-test on time online for noise factor c=1 (double hits). Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.
j10 $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[235] -0.384 (0.701) [258] -9.554 (*) [258] -9.212 (*) [234] -13.262 (*) [235] -12.623 (*) [258] -8.801 (*)
[66] 0.47 (0.712) [258] 0.907 (*) [258] 0.899 (*) [234] 1.0 (*) [235] 0.987 (*) [258] 0.891 (*)
j20 $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[182] -1.204 (0.229) [216] -4.456 (*) [217] -4.913 (*) [177] -11.279 (*) [182] -10.957 (*) [217] -11.457 (*)
[79] 0.43 (0.261) [216] 0.815 (*) [217] 0.825 (*) [177] 0.994 (*) [182] 0.984 (*) [217] 0.968 (*)
j30 $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[182] -2.014 (*) [232] -0.928 (0.353) [234] -2.526 (*) [163] -10.268 (*) [183] -9.31 (*) [233] -9.839 (*)
[119] 0.58 (0.099) [232] 0.69 (*) [234] 0.735 (*) [163] 0.982 (*) [183] 0.94 (*) [233] 0.901 (*)
ubo50 $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[173] -0.722 (0.47) [284] -3.843 (*) [285] -3.933 (*) [162] -11.04 (*) [171] -9.02 (*) [284] -13.107 (*)
[147] 0.524 (0.621) [284] 0.43 (*) [285] 0.435 (*) [162] 1.0 (*) [171] 0.942 (*) [284] 0.951 (*)
ubo100 $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[146] -0.482 (0.63) [288] -8.467 (*) [284] -9.06 (*) [122] -9.585 (*) [141] -6.136 (*) [285] -14.285 (*)
[144] 0.458 (0.359) [288] 0.576 (*) [284] 0.585 (*) [122] 1.0 (*) [141] 0.837 (*) [285] 0.982 (*)
Table 22: Pairwise comparison on time online for noise factor c=2. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each paier method 1 - method 2. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] proportion (p-value) with (*) for p $<$ 0.05
j10 $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[231] 0.391 (0.697) [206] -92.546 (*) [204] -91.532 (*) [233] -4426.055 (*) [230] -4653.225 (*) [205] -1399.699 (*)
$\mathsf{pro}_{0.9}$: 1.01 $\mathsf{pro}_{0.9}$: 0.06 $\mathsf{pro}_{\mathsf{SAA}}$: 0.06 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.03
$\mathsf{pro}_{\mathsf{SAA}}$: 0.99 $\mathsf{stnu}$: 1.94 $\mathsf{stnu}$: 1.94 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.97
j20 $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[173] 0.77 (0.443) [170] -241.897 (*) [172] -262.612 (*) [176] -7787.819 (*) [174] -8817.279 (*) [170] -935.848 (*)
$\mathsf{pro}_{0.9}$: 1.03 $\mathsf{pro}_{0.9}$: 0.03 $\mathsf{pro}_{\mathsf{SAA}}$: 0.03 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.06
$\mathsf{pro}_{\mathsf{SAA}}$: 0.97 $\mathsf{stnu}$: 1.97 $\mathsf{stnu}$: 1.97 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.94
j30 $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[150] -0.264 (0.792) [138] -514.302 (*) [148] -547.413 (*) [160] -14437.976 (*) [152] -14871.238 (*) [140] -397.488 (*)
$\mathsf{pro}_{\mathsf{SAA}}$: 0.99 $\mathsf{pro}_{0.9}$: 0.02 $\mathsf{pro}_{\mathsf{SAA}}$: 0.02 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.07
$\mathsf{pro}_{0.9}$: 1.01 $\mathsf{stnu}$: 1.98 $\mathsf{stnu}$: 1.98 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.93
ubo50 $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{stnu}$-$\mathsf{pro}_{0.9}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[150] -0.707 (0.481) [148] 2600.902 (*) [146] 2585.661 (*) [155] -25795.773 (*) [145] -23210.69 (*) [141] -121.392 (*)
$\mathsf{pro}_{0.9}$: 0.96 $\mathsf{stnu}$: 1.99 $\mathsf{stnu}$: 1.99 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.14
$\mathsf{pro}_{\mathsf{SAA}}$: 1.04 $\mathsf{pro}_{0.9}$: 0.01 $\mathsf{pro}_{\mathsf{SAA}}$: 0.01 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.86
ubo100 $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{pro}_{0.9}$ stnu-$\mathsf{pro}_{0.9}$ $\mathsf{stnu}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{pro}_{\mathsf{SAA}}$-$\mathsf{react}$ $\mathsf{stnu}$-$\mathsf{react}$
[94] 0.141 (0.888) [114] 21743.706 (*) [114] 21911.639 (*) [99] -46429.42 (*) [76] -40238.643 (*) [94] -41.405 (*)
$\mathsf{pro}_{\mathsf{SAA}}$: 1.0 $\mathsf{stnu}$: 2.0 $\mathsf{stnu}$: 2.0 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.21
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.79
Table 23: Magnitude t-test on time online for noise factor c=2 (double hits). Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.